OpenStudy (anonymous):

find the first derivative of y=1-cos2x + 2cos^2x.

OpenStudy (anonymous):

you have to use chain rule with the cos: \[\cos(ax) dx= asin(ax) \]

OpenStudy (anonymous):

ok, thank you!

OpenStudy (anonymous):

np

OpenStudy (dumbcow):

just a note: d/dx cos(x) = -sin(x) so above there should be a neg sign in front of asin(ax)

OpenStudy (anonymous):

ops, confused with integrals hehe. Thanks cow

OpenStudy (anonymous):

\[y=1 - \text{Cos}[2*x] + 2*\text{Cos}[x]^2 = 2 \] Derivative of 2 is zero.